OpenStudy (anonymous):

How can you find the standard form of an equation using 2 given points?

OpenStudy (anonymous):

Calculate your slope and the y intercept.

OpenStudy (anonymous):



OpenStudy (anonymous):

If A is at \((x_{A}, y_{A})\) and B is at \((x_{B}, y_{B})\), then slope is \(\large \frac{y_{B} - y_{A}}{x_{B} - x_{A}}\)

OpenStudy (anonymous):

Ok so what do I do once I have the slope?

OpenStudy (anonymous):

Then you use it to find the y-intercept. Remember, the y-intercept is at point \((0, y_{C})\)

OpenStudy (anonymous):

So, if you used point A, you could use \(\large m = \frac{y_{C} - y_{A}}{0 - x_{A}}\)

OpenStudy (anonymous):

ok so once I have the y-intercept???

OpenStudy (anonymous):

Then put it in the form y = mx + b. m is the slope, and b is the y-intercept.

OpenStudy (anonymous):

Ok,I understand that part but how do i find teh standard form of the equation,my book says it would be Ax +By=C but it doesn texplain how to find the equation using 2 given points.

OpenStudy (anonymous):

Oh, sorry. I missed that part of your post. Standard form is just \(Ax + By = C\) (like you said). You can get it by rearranging your y = mx + b formula.

OpenStudy (anonymous):

Bring y and mx to the same side, to equal b.

OpenStudy (anonymous):

Ok so I have teh x and y intercepts figured out but how do I find C?

OpenStudy (anonymous):

C is just b (from y = mx + b)

OpenStudy (anonymous):

Ok thank you very much!!!

OpenStudy (anonymous):

You're welcome.